# Beyond Win Rates: A Clustering-Based Approach to Character Balance Analysis in Team-Based Games

### v1

@@ 6 Measuring the Impact of Agent Adjustments on Game Balance
7 Limitation and future works
The primary assumption of this study—that agents with similar team compositions are perfect substitutes—presents several limitations. While this assumption provides a useful starting point for analyzing agent relationships, it oversimplifies the complexities of team compositions and gameplay dynamics. In reality, agent roles are not strictly interchangeable even when agents share similar co-occurrence patterns in team compositions. The strategic context in which agents are chosen can vary significantly, and players may select agents based not only on their abilities but also on personal playstyle preferences, team synergy, and broader game strategy. For instance, two agents with similar co-occurrence data may still play distinct roles depending on the player’s preferred tactics, such as an agent focusing on aggressive plays versus one providing support and utility. Therefore, assuming perfect substitutability risks overlooking these strategic nuances, which can have a significant impact on gameplay balance.
Moreover, while clustering agents based on team composition provides valuable insights, the exact definition of "balance" remains elusive. The distance metric used in this study captures how similar agents are in terms of team composition, but it does not inherently provide a clear boundary between balanced and imbalanced agents. In other words, the clustering approach does not establish a direct link between the distance measure and game balance—meaning we cannot definitively conclude whether an agent’s distance from others indicates that it is "too strong" or "too weak" in the context of game balance. This absence of a clear threshold for balance presents a challenge for developers seeking to fine-tune agent strength or adjust their roles within team compositions.
Future work could explore incorporating additional factors, such as side-specific win rates, into the clustering process. Including this metric would allow for a deeper understanding of how agent compositions perform under different conditions and could provide a more comprehensive view of balance. Additionally, incorporating more granular factors like player behavior or contextual gameplay data could refine this approach. Developing robust metrics for determining "balance" based on distance and adding subjective elements, such as strategy and player preference, would further enhance the utility of this method in real-world game balancing, enabling a more nuanced and flexible assessment of agent roles and team compositions.

### v2

@@ VII Measuring the Impact of Agent Adjustments on Game Balance
VIII Conclusion
Traditional balance metrics offer limited insight into team-based game dynamics. This paper proposed and validated a novel clustering methodology for character balance in Valorant using professional match data and team composition tendencies. Applying HAC with JSD to agent co-occurrence identified distinct agent clusters aligning with established roles and highlighting unique tactical profiles and nuanced synergies beyond predefined categories.
JSD distances also provide a quantitative framework to assess balance adjustment impacts on agent roles and team dynamics, enabling more data-driven, iterative game design. The primary contribution is a methodology offering a holistic, interpretable, context-aware perspective on character balance. By focusing on collective agent function, our approach complements existing metrics, providing developers a tool for informed adjustments promoting strategic diversity. Despite limitations, this method represents a significant step towards a more nuanced understanding of game balance.
IX Limitation and Future Works
Our primary assumption—that agents with similar team compositions are perfect substitutes—oversimplifies the complexities of gameplay. Agent roles are not strictly interchangeable, as players choose agents based on playstyle, synergy, and broader strategy. Two agents with similar co-occurrence data may still play distinct roles. The clustering reveals similarity in team composition, but the distance metric alone doesn’t define a ”balanced” vs. ”imbalanced” threshold. This poses a challenge for translating the analysis into specific balance adjustments.
Future work could incorporate factors like side-specific win rates or player behavior for a more comprehensive view of balance. At the same time, developing robust metrics for determining ”balance” based on distance is crucial. While the impact of specific character skill changes on win rates varies greatly between different games and even different patches within a single game, developers can establish a relative sense of ”overly large” changes. This could be achieved by normalizing the magnitude of JSD changes observed after a patch against a historical baseline of changes from previous patches. Furthermore, correlating these quantitative JSD shifts with qualitative community feedback (e.g., forum discussions, surveys) could provide a valuable cross-validation, helping to identify changes that are perceived as disruptive by the player base, even if their direct impact on win rates is difficult to isolate. Adding subjective elements like strategy would enhance the method’s utility for real-world game balancing.